Profitability Ratios

Financial statements for Remington Inc. follow.

Remington Inc.
Consolidated Statements of Income
(In thousands except per share amounts)
2019 2018 2017
Net sales $7,245,088 $6,944,296 $6,149,218
Cost of goods sold (5,286,253) (4,953,556) (4,355,675)
Gross margin $1,958,835 $1,990,740 $1,793,543
General and administrative expenses (1,259,896) (1,202,042) (1,080,843)
Special and nonrecurring items 2,617 - -
Operating income $701,556 $788,698 $712,700
Interest expense (63,685) (62,398) (63,927)
Other income 7,308 10,080 11,529
Gain on sale of investments - 9,117 -
Income before income taxes $645,179 $745,497 $660,302
Provision for income taxes 254,000 290,000 257,000
Net income $391,179 $455,497 $403,302
Net income per share $1.08 $1.25 $1.11
Remington Inc.
Consolidated Balance Sheets
(In thousands)
ASSETS Dec. 31, 2019 Dec. 31, 2018
Current assets:
Cash and equivalents $320,558 $41,235
Accounts receivable 1,056,911 837,377
Inventories 733,700 803,707
Other 109,456 101,811
Total current assets $2,220,625 $1,784,130
Property and equipment, net 1,666,588 1,813,948
Other assets 237,392 248,372
Total assets $4,124,605 $3,846,450
LIABILITIES AND STOCKHOLDERS' EQUITY
Current liabilities:
Accounts payable $250,363 $309,092
Accrued expenses 347,892 274,220
Other current liabilities 15,700 -
Income taxes 93,489 137,466
Total current liabilities $707,444 $720,778
Long-term debt $650,000 $541,639
Deferred income taxes 275,101 274,844
Other long-term liabilities 61,267 41,572
Total liabilities $1,693,812 $1,578,833
Stockholders' equity:
Common and preferred stock $189,727 $189,727
Additional paid-in capital 128,906 127,776
Retained earnings 2,397,112 2,136,794
$2,715,745 $2,454,297
Less: Treasury stock, at cost (284,952) (186,680)
Total stockholders' equity $2,430,793 $2,267,617
Total liabilities and stockholders' equity $4,124,605 $3,846,450

Use the information provided above and below to respond to the following requirements.

Statement Item January 1, 2018 (in thousands)
Total assets $3,485,233
Total stockholders’ equity 2,083,122

Required:

Compute the five profitability ratios for 2018 and 2019. Round intermediate calculations and final answers to two decimal places.

2019 2018
Gross profit percentage % %
Operating margin percentage % %
Net profit margin percentage % %
Return on assets % %
Return on equity % %
Answer #1
1)
Gross Margin Percentage
2019 2018
= Gross Margin/ Sales 27.04% 28.67%
=1958835/7245088 =1990740/6944296
2)
Operating Margin Percentage 2019 2018
9.68% 11.36%
=Operating Income/ Sales =701556/7245088 =788698/6944296
3)
Net Profit Margin Percentage 2019 2018
5.40% 6.56%
= Net Income/ Sales =391179/7245088 =455497/6944296
4)
Return On Assets 2019 2018
9.81% 12.43%
= Net Income / Average Operating Assets
=391179/(4124605+3846450)/2 =455497/(3485233+3846450)/2
5) 2019 2018
Return On Equity 16.65% 20.94%
= Net Income / Average Shareholders Equity =391179/(2430793+2267617)/2 =455497/(2267617+2083122)/2
Feel Free To Discuss Queries.          Please Rate